Using the data from your Table 2, fill the following table Table 3 - Calculating the period of oscillation of a pendulum Period AT Trial

Data from table 2

Part 2 - Time measurement [1] Table 2 - Calculating the period of oscillation of a pendulum Time for 10 oscillations Trial At +o 35 to. 35 21247 ra 4 o, 3 Laboratory report: Simple measurements & buoyancy force
Answer #1

The table gives the data for the time taken to complete 10 oscillations. Time period (or period of oscillation) is the time it takes to complete 1 oscillation. So, period of oscillation here will be:

T = t/10

for example, take trial 1,

t = 13.25s

therefore, period of oscillation for trial 1 is: T = 13.25/10 = 1.325s

\Delta T is the uncertainty in the period of oscillation and so it is equal to the measured uncertainty of the instrument.

\Delta T = \Delta t

so, for trial 1, \Delta T = \pm 0.35s

repeat for other trials.